### v1

# Inner products of deep level Deligne–Lusztig representations of Coxeter type

Abstract.
In this article we prove orthogonality relations for deep level Deligne–Lusztig varieties of Coxeter type, attached to a reductive group over a local non-archimedean field, which splits over an unramified extension. This extends results of [DI24], where the quasi-split case was handled. This allows to construct new irreducible representations of parahoric subgroups of $p$-adic groups.
1. Introduction
Let $k$ be a non-Archimedean local field with residue characteristic $p>0$, integers $\mathcal{O}_{k}$, uniformizer $\varpi$ and residue field $\mathbb{F}_{q}$. Let $\breve{k}$ be the completion of the maximal unramified extension of $k$, let $\mathcal{O}_{{\breve{k}}}$ denote the integers of $\breve{k}$. Let $F$ denote the Frobenius automorphism of $\breve{k}$ over $k$.
Let $G$ be a reductive group over $k$, which splits over $\breve{k}$. Let $T\subseteq B$ be a maximal torus and a Borel subgroup of $G$, such that $T$ splits and $B$ becomes rational over $\breve{k}$. Denote by $U$ the unipotent radical of $B$ and assume that $(T,U)$ is a Coxeter pair (see §2.1). In particular, $T$ is elliptic and the apartment of $T$ in the reduced Bruhat–Tits building of $G$ consists of one point. Bruhat–Tits theory attaches to this point a (connected) parahoric $\mathcal{O}_{k}$-model $\mathcal{G}$ of $G$. For any $r\leq\infty$ we can regard $\mathcal{G}(\mathcal{O}_{{\breve{k}}}/\varpi^{r})=\mathbb{G}_{r}(\overline{}\mathbb{F}_{q})$ as the geometric points of a perfect ¯$\overline{}\mathbb{F}_{q}$-scheme $\mathbb{G}_{r}$. This is done via the (truncated, if $r<\infty$) positive loop functor, see e.g. [Zhu17, §1.1] (or [DI24, §2]) for details.
We fix now some $r<\infty$ and write $\mathbb{G}=\mathbb{G}_{r}$. If $\mathcal{G}$ is defined over $\mathcal{O}_{k}$, then $\mathbb{G}$ is defined over $\mathbb{F}_{q}$; in this case we also denote by $F$ the geometric Frobenius of $\mathbb{G}$, so that $\mathbb{G}^{F}=\mathbb{G}(\mathbb{F}_{q})$. Moreover, if $H\subseteq G$ is a subscheme, then we denote by $\mathcal{H}$ its closure in $\mathcal{G}$ and by $\mathbb{H}\subseteq\mathbb{G}$ the corresponding subscheme of $\mathbb{G}$. Consider the closed subscheme
(1.1) $X=X^{\mathcal{G}}_{T,U,r}=\{x\in\mathbb{G}\colon x^{-1}F(x)\in F\mathbb{U}\}$
of $\mathbb{G}$. The group $\mathbb{G}^{F}\times\mathbb{T}^{F}$ acts on it by $(g,t)\colon xgxt$. Let $\ell\neq p$ be a prime. For any smooth character $\chi\colon\mathbb{T}^{F}\rightarrow\overline{}\mathbb{Q}_{\ell}^{\times}$ we have the virtual $\mathbb{G}^{F}$-module
$H_{c}^{\ast}(X)[\chi]=\sum_{i\in\mathbb{Z}}(-1)^{i}H^{i}(X_{T,U},\overline{}\mathbb{Q}_{\ell}^{\times})[\chi],$
where for any $\mathbb{T}^{F}$-module $M$, $M[\chi]$ denotes the $\chi$-isotypic subspace. The following is our main result.
Theorem 1.1.
Suppose that $q$ satisfies condition (2.1) (this is always true when $q>5$). Then there exists a Coxeter pair $(T,U)$ such that
$\dim_{\overline{}\mathbb{Q}_{\ell}}\operatorname{Hom}_{\mathbb{G}^{F}}(H_{c}^{\ast}(X)[\chi],H_{c}^{\ast}(X)[\chi^{\prime}])=\sharp\left\{w\in W_{e}^{F};w(\chi)=\chi^{\prime}\right\}$
for any two smooth characters $\chi,\chi^{\prime}$ of $\mathbb{T}^{F}$, where $W_{e}$ denotes the Weyl group of the special fiber of $\mathcal{T}$ in the reductive quotient of the special fiber of $\mathcal{G}$.
Remark 1.2.
Recently, under a mild condition on $p$, Chan [Cha24] shows that the inner product formula in Theorem 1.1 holds if $(T,\theta)$ is split-generic, using a different approach.
In particular, if $\{w\in W_{e}^{F}\colon w(\chi)=\chi\}=\{1\}$, then $H_{c}^{\ast}(X)[\chi]$ is up to sign an irreducible $\mathbb{G}^{F}$-representation. Note that Theorem 1.1 generalizes [DI24, Theorem 3.2.3] and [CI23, Theorem 4.1].
Now we comment on Theorem 1.1. First, we explain why “it suffices” to establish the theorem for a single Coxeter pair $(T,U)$. Ultimately, we are interested in smooth representations of the $p$-adic group $G(k)$. One has the $p$-adic Deligne–Lusztig space $X_{T,U}\cong\dot{X}_{\dot{c}}(b)$ from [Iva23a, §7.2 and §11.2] equipped with an action of $G(k)\times T(k)$. By [Nie24, Iva23b], one has
$X_{T,U}=\coprod_{\gamma\in G(k)/\mathcal{G}(\mathcal{O}_{k})}\gamma X_{T,U,\infty}^{\mathcal{G}}.$
Thus for any smooth character $\theta\colon T(k)\rightarrow\overline{}\mathbb{Q}_{\ell}^{\times}$, the $\theta$-isotypic part of the cohomology of $X_{T,U}$ equals
$R_{T,U}(\theta)={\text{c-}\rm Ind}_{\mathcal{G}(\mathcal{O}_{k})Z(k)}^{G(k)}H^{\ast}_{c}(X)[\theta|_{\mathcal{T}(\mathcal{O}_{k})}],$
where $Z\subseteq G$ is the center and $H^{\ast}_{c}(X)[\theta|_{\mathcal{T}(\mathcal{O}_{k})}]$ is extended to a representation of $\mathcal{G}(\mathcal{O}_{k})Z(k)$ in the unique way such that its central character is $\theta|_{Z(k)}$. Now, the point is that by [Iva23a, Corollary 7.25, Lemma 11.3], $X_{T,U}$ are all mutually $G(k)\times T(k)$-equivariantly isomorphic, when $(T,U)$ varies through all Coxeter pairs $(T,U)$ with a fixed $T$. Thus, the $G(k)$-representation $R_{T,U}(\theta)$ is independent of the choice of $U$, as long as $(T,U)$ remains Coxeter. So, it suffices to know the statement of the theorem for at least one Coxeter pair. In fact, our proof shows that for many groups $G$ Theorem 1.1 holds for all pairs $(T,U)$, see Remark 2.4.
Next, we explain why the condition on $q$ in the theorem is very mild, so that Theorem 1.1 even allows to construct new irreducible representations. Recall that by the work of Yu and Kaletha [Yu01, Kal19], one can attach a supercuspidal irreducible $G(k)$-representation $\pi_{(S,\theta)}$ to any regular elliptic pair $(S,\theta)$ consisting of a maximal elliptic torus $S\subseteq G$ and a sufficiently nice smooth character $\theta\colon S(k)\rightarrow\overline{}\mathbb{Q}_{\ell}^{\times}$. A crucial point for this to work is the existence of a Howe factorization of $\theta$, cf. [Kal19, §3.6]. However, not all characters admit a Howe factorization, when the residue characteristic $p$ is small and $G$ is not an inner form of $\mathrm{GL}_{n}$.
For instance, if $p\in\{2,3,5\}$, there exist many examples of pairs $(T,\theta)$ with $T$ unramified Coxeter (hence covered by our main result when $q$ satisfied condition (2.1) – in particular, whenever $q>5$) such that ${\rm Stab}_{W_{e}^{F}}(\theta)=\{1\}$, but $\theta$ does not admit a Howe factorization. For examples of $(T,\theta)$ not admitting a Howe factorization we refer to the forthcoming work of Fintzen–Schwein [FS], where an algebraic approach to the extension of Yu’s construction is pursued. As ${\rm Stab}_{W_{e}^{F}}(\theta)=\{1\}$, one should expect an irreducible supercuspidal $G(k)$-representation attached to $(T,\theta)$, but Yu’s construction does not apply as there is no Howe factorization. The point is now that our cohomological construction does not require any condition on $p$, but only a mild one on $q$. In particular, there are many examples of $k,G,T,\theta$ such that ±$\pm H^{\ast}_{c}(X)[\theta|_{\mathcal{T}(\mathcal{O}_{k})}]$ is an irreducible $\mathcal{G}(\mathcal{O}_{k})$-representation, which does not appear in Yu’s construction.
However, let us also note that our method is not yet complete, in the sense that it does not prove that the resulting $G(k)$-representation $R_{T,U}(\theta)$ is irreducible supercuspidal. So far there is no purely Deligne–Lusztig theoretic proof of this fact; the closest purely Deligne–Lusztig theoretic result states (in the case of inner forms of $\mathrm{GL}_{n}$) that $R_{T,U}(\theta)$ is admissible and hence a direct sum of finitely many irreducible supercuspidal representations, see [CI23, Theorem 6.1].
Acknowledgements
The first author is grateful to Jessica Fintzen and David Schwein for explaining their results on characters without Howe decomposition. The first author gratefully acknowledges the support of the German Research Foundation (DFG) via the Heisenberg program (grant nr. 462505253).
2. Preparations
2.1. More notation
We use the notation from the introduction. Moreover, we denote by $Z$ the center of $G$, $N_{G}(T)$ the normalizer of $T_{\breve{k}}$ in $G_{\breve{k}}$, by $W=N_{G}(T)/T$ the Weyl group of $T$, by $X^{\ast}(T)$ (resp. $X_{\ast}(T)$) the group of characters (resp. cocharacters) of $T_{\breve{k}}$ and by $\langle\cdot,\cdot\rangle\colon X^{\ast}(T)\times X_{\ast}(T)\mathbb{Z}$ the natural pairing. We write $\Phi$ for the root system of $T_{\breve{k}}$ in $G_{\breve{k}}$, $\Phi^{+}$ for the subset of positive roots determined by $B$, and $\Delta\subseteq\Phi^{+}$ for the subset of positive simple roots. We write $S\subseteq W$ for the corresponding set of simple reflections.
Let $c\in W$ be the unique element such that $FB={}^{c}B$. Then for any lift $\dot{c}$ of $c$, ${\mathrm{Ad}}({\dot{c}})^{-1}\circ F:G(\breve{k})G(\breve{k})$ fixes the pinning $(T,B)$, hence defines automorphisms, denoted by $\sigma$, of the based root system $\Delta\subseteq\Phi$ and of the Coxeter system $(W,S)$. Note that $\sigma$ does not depend on the choice of the lift $\dot{c}$. We call $(T,B)$ (or $(T,U)$) a Coxeter pair if $c$ is a $\sigma$-Coxeter element in the Coxeter triple $(W,S,\sigma)$, that is, if a(ny) reduced expression of $c$ contains precisely one element from each $\sigma$-orbit on $S$. Moreover, we assume throughout the article that $c$ is $\sigma$-Coxeter, and hence $(T,U)$ is a Coxeter pair.
Except for $G$, $\mathcal{G}$ and their subgroups (which are defined over $k,\breve{k}$ resp. $\mathcal{O}_{k},\mathcal{O}_{\breve{k}}$), all schemes appearing below are perfect schemes perfectly of finite presentation and perfectly smooth over ¯$\overline{}\mathbb{F}_{q}$. For a review of perfect geometry we refer to [Zhu17, Appendix A]. We freely make use of the $6$-functor formalism of étale cohomology for such schemes with $\overline{}\mathbb{Q}_{\ell}$-coefficients. Moreover, for a perfect ¯$\overline{}\mathbb{F}_{q}$-scheme we denote by $H_{\ast}(Y)=H_{c}^{\ast}(Y,\overline{}\mathbb{Q}_{\ell})$ its $\ell$-adic étale cohomology with compact support.
2.2. Pinning
We may express the action of the Frobenius $F$ on $X_{\ast}(T)_{\mathbb{Q}}$ as $F=\mu c\sigma\colon x\mu+c\sigma(x)$ for some $\mu\in X_{\ast}(T)$. There is a unique point $e\in\mathbb{Q}\Phi^{\vee}$ such that $F(e)\in e+X_{\ast}(Z)_{\mathbb{Q}}$, or equivalently, $\mu+c\sigma(e)-e\in X_{\ast}(Z)_{\mathbb{Q}}$. Let
$\Phi_{e}=\{\alpha\in\Phi;\langle\alpha,e\rangle\in\mathbb{Z}\}.$
We denote by $\Delta_{e}$ the set of simple roots of $\Phi_{e}^{+}=\Phi_{e}\cap\Phi^{+}$. Let $W_{e}\subseteq W$ be the Weyl group of $\Phi_{e}$. Note that $\mathcal{G}$ from the introduction is the parahoric model attached to the image of $e$ in the reduced building of $G$, and that $\Phi_{e}$ (resp. $W_{e}$) is the root system (resp. Weyl group) of the reductive quotient of the special fiber of $\mathcal{G}$.
Also, note that the action of $F$ on $W$ agrees with ${\rm Ad}(c)\circ\sigma$; we denote it by $F=c\sigma\colon WW$. This action stabilizes $W_{e}\subseteq W$. Finally, for an element $w\in W_{e}$ we denote by $\dot{w}\in\mathbb{G}(\overline{}\mathbb{F}_{q})$ an arbitrary (fixed) lift of $w$.
2.3. A condition on $q$
Let $\omega_{\alpha}$ denotes the fundamental coweight of $\alpha\in\Delta$. For a $\sigma$-orbit $\mathcal{O}\subseteq\Delta$ of simple roots, we set $\omega_{\mathcal{O}}^{\vee}=\sum_{\alpha\in\mathcal{O}}\omega_{\alpha}^{\vee}$, where $\omega_{\alpha}^{\vee}$ denotes the fundamental coweight of $\alpha\in\Delta$. We prove our main result under the following condition on $q$:
(2.1) $q>M=\max\{\langle\gamma,\omega_{\mathcal{O}}^{\vee}\rangle;\gamma\in\Phi^{+},\mathcal{O}\in\Delta/\langle\sigma\rangle\}.$
Note that $M$ only depends on the (relative) Dynkin diagram $\Delta$ of the quasi-split inner form of $G$ over $k$. If $\Delta$ is connected then $M$ takes the following values: $M=1$ for type $A_{n}$; $M=2$ for types $B_{n},C_{n},D_{n},{}^{2}A_{n},{}^{2}D_{n}$; $M=3$ for types $G_{2},E_{6},{}^{3}D_{4}$; $M=4$ for types $F_{4},E_{7},{}^{2}E_{6}$; $M=6$ for type $E_{8}$. If the quasi-split inner form of $G$ is split, then $M$ is the same as in [DI24, §2.7], and it differs otherwise. Just as in [DI24, §2.7], for arbitrary $G$ the constant $M$ equals the maximum of the values of $M$ over all connected components of the Dynkin diagram of $G_{\breve{k}}$ (equipped with the smallest power of $\sigma$ fixing the connected component). In particular, (2.1) holds whenever $q>5$.
2.4. A Coxeter element in $W_{e}$
It turns out that $c$ determines a (twisted) Coxeter element of $W_{e}$. Write $c=s_{\alpha_{1}}\cdots s_{\alpha_{r}}$, where $\{\alpha_{1},\dots,\alpha_{r}\}\subseteq\Delta$ is a set of representatives of $\sigma$-orbits of $\Delta$.
Let $I=(i_{1}<i_{2}<\cdots<i_{m})$ be a subsequence of $[r]:=(1<2<\cdots<r)$, and let $I^{\prime}=(j_{1}<j_{r}<\cdots<j_{r-m})$ be the complement sequence of $I$ in $[r]$. We define
$\sigma_{I,c}$ $=s_{\alpha_{i_{1}}}s_{\alpha_{i_{2}}}\cdots s_{\alpha_{i_{m}}}\sigma;$
$c_{I}$ $=s_{\beta_{j_{1}}}s_{\beta_{j_{2}}}\cdots s_{\beta_{j_{r-m}}};$
$\Delta_{I,c}$ =$=\{\beta_{j_{l}};1\leqslant l\leqslant r-m\}$
where $\beta_{j_{l}}=s_{\alpha_{i_{1}}}s_{\alpha_{i_{2}}}\cdots s_{\alpha_{i_{t}}}(\alpha_{j_{l}})$ with $11\leqslant t\leqslant m-$ 1 such that $i_{t}<j_{l}<i_{t+1}$. By definition, $c\sigma=c_{I}\sigma_{I,c}$.
Theorem 2.1.
Let $c$, $\mu$ and $e=e_{\mu,c}$ be as in §2.2. Then there exist a sequence $I=I_{\mu,c}$ of $1<2<\cdots<r$ such that
(1) $\sigma_{I,c}(\Delta_{e})=\Delta_{e}$;
(2) $\Delta_{I,c}\subseteq\Delta_{e}$ is a representative set of $\sigma_{I,c}$-orbits of $\Delta_{e}$;
(3) $\sigma_{I,c}^{i}=1$ if and only if $\sigma_{I,c}^{i}$ fixes each root of $\Delta_{e}$.
In particular, $c_{I}$ is a $\sigma_{I,c}$-Coxeter element of $W_{e}$.
This theorem is proven in §4.
2.5. Support
For $\alpha\in\Phi$ we denote by $\operatorname{supp}(\alpha)\subseteq\Delta$ the minimal subset whose linear span contains $\alpha$. For a subset $C\subseteq\Phi$ we set $\operatorname{supp}(C)=\cup_{\alpha\in C}\operatorname{supp}(\alpha)$. For $w\in W$ we denote by $\operatorname{supp}(w)$ the set of simple reflections which appear in some/any reduced expression of $w$.
Lemma 2.2.
Let $C\subseteq\Phi$ be a $c\sigma$-orbit. Then $\operatorname{supp}(C)$ is $\sigma$-stable.
Proof.
Let $c=s_{\alpha_{1}}\cdots s_{\alpha_{r}}$ be as in §2.4. Let $\alpha\in\operatorname{supp}(\gamma)$ for some $\gamma\in C$. It suffices to show that the $\sigma$-orbit $\mathcal{O}$ of $\alpha$ is contained in $\operatorname{supp}(C)$. Set $\delta=\sharp\mathcal{O}$. Let $1\leqslant j\leqslant r$ be the unique integer such that $\alpha_{j}\in\mathcal{O}$. Let $10\leqslant i_{0}\leqslant\delta-$ 1 such that
$\alpha,\sigma^{-1}(\alpha),\dots,\sigma^{1-i_{0}}(\alpha)\neq\alpha_{j}\text{ and }\sigma^{-i_{0}}(\alpha)=\alpha_{j}.$
Then one checks that $(c\sigma)^{-i_{0}}=\sigma^{-i_{0}}w$ for some $w\in W$ such that $\operatorname{supp}(w)\in S-\{s_{\alpha}\}$. Hence $\alpha\in\operatorname{supp}(w(\gamma))$ and $\alpha_{j}=\sigma^{-i_{0}}(\alpha)\in\operatorname{supp}(\sigma^{-i_{0}}w(\gamma))=\operatorname{supp}((c\sigma)^{-i_{0}}(\gamma))$. So we can assume that $\alpha=\alpha_{j}$. Let $10\leqslant i\leqslant\delta-$ 1. Note that $(c\sigma)^{i}=u_{i}\sigma^{i}$ for some $u_{i}\in W$ with $\operatorname{supp}(u_{i})\in S-\{\sigma^{i}(\alpha_{j})\}$. It follows that $\sigma^{i}(\alpha)\in\operatorname{supp}((c\sigma)^{i}(\gamma))$. So the statement follows. ∎
Proposition 2.3.
Let $C$ be a $c\sigma$-orbit of $\Phi$. Then $\operatorname{supp}(C)=\cup_{i\in\mathbb{Z}}\sigma^{i}(H)$, where $H$ is a connected component of $\Delta$.
Proof.
Without loss of generality we may assume that $\Delta=\cup_{i\in\mathbb{Z}}\sigma^{i}(H)$. We argue by induction on $\sharp\Delta$. Assume the statement is false. Let $c=s_{\alpha_{1}}\cdots s_{\alpha_{r}}$ be as in §2.4. By Lemma 2.2 there exists $1\leqslant j\leqslant r$ such that $C\subseteq\Phi_{K}$, where $K=\Delta-\mathcal{O}$ and $\mathcal{O}$ is the $\sigma$-orbit of $\alpha_{j}$. By replacing $c$ with its $W_{K}$-$\sigma$-conjugate $s_{\alpha_{j}}\cdots s_{\alpha_{r}}\sigma(s_{\alpha_{1}}\cdots s_{\alpha_{j-1}})$, we can assume that $j=1$ and $\alpha_{1}\in\mathcal{O}$. Let $c^{\prime}=s_{\alpha_{1}}c$, which is a $\sigma$-Coxeter element of $W_{K}$. As $C\subseteq\Phi_{K}$, $C$ is also a $c^{\prime}\sigma$-orbit of $\Phi_{K}$. By induction hypothesis we have $\operatorname{supp}(C)=\cup_{i\in\mathbb{Z}}\sigma^{i}(D)$, where $D$ is a connected component of $H-\{\alpha_{1}\}$. As $H$ is connected, there exists $\gamma\in C$ and $\beta\in\operatorname{supp}(\gamma)$ such that
$0>\langle\alpha_{1},\beta^{\vee}\rangle\geqslant\langle\alpha_{1},\gamma^{\vee}\rangle.$
Then we have $\sigma^{-1}(\alpha_{1})\in\operatorname{supp}((c\sigma)^{-1}(\gamma))$, contradicting that $C\subseteq\Phi_{K}$. The proof is finished. ∎
2.6. A condition on the $\sigma$-Coxeter element
Let $c$, $\mu$, $e=e_{\mu,c}$, $I=I_{\mu,c}$, $c_{I}$, $\sigma_{I}=\sigma_{I,c}$ and $\Delta_{I}=\Delta_{I,c}\subseteq\Delta_{e}$ be as in Theorem 2.1. Denote by $\ell\colon W\mathbb{Z}_{\geq 0}$ (resp. $\ell_{e}:W_{e}\mathbb{Z}_{\geqslant 0}$) the length function associated to the set $\Delta$ (resp. $\Delta_{e}$) of simple roots. Let $w_{0}$ and $w_{e}$ be the longest elements of $W$ and $W_{e}$ respectively. We consider the following condition on $c$, or, equivalently, on the pair $(T,U)$:
($\ast$) $\text{There exists $N\in\mathbb{Z}_{\geqslant 1}$ such that }\,(c\sigma)^{N}=w_{0}\sigma^{N},\quad N\ell(c)=\ell(w_{0}).$
Remark 2.4.
If $\Delta$ is connected, then there always exists a $\sigma$-Coxeter element $c\in W$ satisfying ($\ast$2.6), see [Bou68, Chap. V, Prop. 6.2]. Moreover, if the Coxeter number of $G$ is even, then any $c$ satisfies this condition.
Lemma 2.5.
Suppose $c$ satisfies condition (italic-$\ast$2.6). Then $(c_{I}\sigma_{I})^{N}=w_{e}\sigma_{I}^{N}$ and $N\ell_{e}(c_{I})=\ell_{e}(w_{e})$.
Proof.
By Theorem 2.1, $c_{I}\sigma_{I}=c\sigma$ and $\sigma_{I}(\Delta_{e})=\Delta_{e}$. As $(c_{I}\sigma_{I})^{N}=w_{0}\sigma^{N}$, it follows that $(c_{I}\sigma_{I})^{N}$ sends $\Phi_{e}^{+}$ to $-\Phi_{e}^{+}$, that is, $(c_{I}\sigma_{I})^{N}=w_{e}\sigma_{I}^{N}$.
It remains to show $\ell_{e}((c_{I}\sigma_{I})^{i+1})=\ell_{e}((c_{I}\sigma_{I})^{i})+\ell_{e}(c_{I}\sigma_{I})$ for $11\leqslant i\leqslant N-$ 1. Indeed, this is equivalent to that for any $\alpha\in\Phi_{e}^{+}$ with $(c_{I}\sigma_{I})^{-1}(\alpha)<0$ we have $(c_{I}\sigma_{I})^{i}(\alpha)>0$. This statement follows from that $c_{I}\sigma_{I}=c\sigma$ and $\ell((c\sigma)^{i+1})=\ell((c\sigma)^{i})+\ell(c\sigma)$ for $11\leqslant i\leqslant N-$ 1. ∎
For $w\in W$ we denote by $\operatorname{supp}(w)$ the set of simple reflections in $\Delta$ that appears in some/any reduced expression of $w$. For $u\in W_{e}$, we can define $\operatorname{supp}_{\Delta_{e}}(u)\subseteq\Delta_{e}$ in a similar way.
Corollary 2.6.
Suppose $c$ satisfies condition (italic-$\ast$2.6). Let $K\subsetneq\Delta_{e}$ be a proper $\sigma_{I}$-stable subset. Then there exists a proper $\sigma$-stable subset $J\subsetneq\Delta$ such that $\sigma_{I}\in W_{J}\sigma$ and $w_{e}W_{K}\subseteq w_{0}W_{J}$.
Proof.
Let notation be as in §2.4. As $\Delta_{I}=\{\beta_{j};j\in I^{\prime}\}$ with $I^{\prime}=[r]-I$ is a representative set of $\Delta_{e}$, there exists $i\in I^{\prime}$ such that $\beta_{i}\notin K$. Let $J=\Delta-\mathcal{O}_{i}$, where $\mathcal{O}_{i}$ is the $\sigma$-orbit of $\alpha_{i}$. By construction, $\operatorname{supp}(s)\subseteq J$ for $s\in K$ and $\operatorname{supp}(\sigma_{I}\sigma^{-1})\subseteq J$. By Lemma 2.5 we have
$w_{e}=(c_{I}\sigma_{I})^{N}\sigma_{I}^{-N}=(c\sigma)^{N}\sigma_{I}^{-N}=w_{0}\sigma^{N}\sigma_{I}^{-N}\subseteq w_{0}W_{J}.$
Thus $w_{e}W_{K}\subseteq w_{0}W_{J}$ as desired. ∎
Lemma 2.7.
Let $K_{1},K_{2}\subseteq\Delta_{e}$ be two $\sigma_{I}$-stable subsets. Let $c_{1}$ and $c_{2}$ be two $\sigma_{I}$-Coxeter elements of $W_{K_{1}}$ and $W_{K_{2}}$ respectively. Let $w\in W_{e}$ such that $c_{1}\sigma_{I}(w)=wc_{2}$. Then there exists $x\in{}^{K_{1}}W_{e}{}^{K_{2}}$ such that ${}^{x}K_{2}=K_{1}$ and $w\in xW_{K_{2}}$.
Proof.
By symmetry we may assume $\sharp K_{1}\leqslant\sharp K_{2}$. Let $x\in{}^{K_{1}}W_{e}$ such that $w\in W_{K_{1}}x$. Then there exists $c_{2}^{\prime}\leq c_{2}$ such that $xc_{2}\in W_{K_{1}}xc_{2}^{\prime}$ and $xc_{2}^{\prime}\in{}^{K_{1}}W_{e}$. Hence we have $\sigma_{I}(x)=xc_{2}^{\prime}$. Note that $c_{2}^{\prime}$ is a partial $\sigma_{I}$-Coxeter element, which is of minimal length (in the sense of $\ell_{e}$) in its $\sigma_{I}$-conjugacy class. Thus $c_{2}^{\prime}=1$, $x=\sigma_{I}(x)$ and $x(\operatorname{supp}_{\Delta_{e}}(c_{2}))\subseteq K_{1}$, which implies that $x(K_{2})\subseteq K_{1}$. Hence $x(K_{2})=K_{1}$ since $\sharp K_{1}\leq\sharp K_{2}$. Thus $x\in{}^{K_{1}}W_{e}{}^{K_{2}}$ as desired. ∎
3. Cohomology of $X$
Recall the scheme $X$ from (1.1) equipped with $\mathbb{G}^{F}\times\mathbb{T}^{F}$-action.
3.1. The schemes $\Sigma^{i}$
Let $i\in\mathbb{Z}$. We define
$\Sigma^{i}=\{(x,x^{\prime},y)\in\mathbb{U}\times F^{i+1}\mathbb{U}\times\mathbb{G};xF(y)=yx^{\prime}\}.$
Let ${\rm pr}_{3}:\Sigma^{i}\mathbb{G}$ be the natural projection. There is a locally closed decomposition
$\Sigma^{i}=\bigsqcup_{w\in W_{e}}\Sigma_{w}^{i},$
where $\Sigma_{w}^{i}={\rm pr}_{3}^{-1}(\mathbb{U}w\mathbb{T}\mathbb{G}^{1}F^{i}\mathbb{U})$.
The group $\mathbb{T}^{F}\times\mathbb{T}^{F}$ acts on $\Sigma^{i}$ and on each of the pieces $\Sigma^{i}_{w}$ by
$(t,t^{\prime}):(x,x^{\prime},y)(txt^{-1},t^{\prime}x^{\prime}{t^{\prime}}^{-1},ty{t^{\prime}}^{-1}).$
As in [DL76, p.137] there is a $\mathbb{T}^{F}\times\mathbb{T}^{F}$-equivariant isomorphism $X\times X/\mathbb{G}^{F}\stackrel{{\scriptstyle\sim}}{{}}\Sigma^{0}$, and for characters $\chi,\chi^{\prime}$ of $\mathbb{T}^{F}$ we have
$\dim_{\overline{\mathbb{Q}}_{\ell}}\operatorname{Hom}_{\mathbb{G}^{F}}(H_{*}(X)[\chi^{\prime}],H_{*}(X)[\chi])=\dim H_{*}(\Sigma^{0})_{\chi^{\prime},\chi^{-1}},$
where $H_{*}(\Sigma^{0})_{\chi^{\prime},\chi}$ is the corresponding isotropic subspace of $H_{c}^{*}(\Sigma^{0})$.
Let $Z\subseteq G$ denote the centre of $G$ and consider the embedding $z(z,z^{-1})\colon ZT\times T$. Then the above $\mathbb{T}^{F}\times\mathbb{T}^{F}$-action on $\Sigma^{i}$ factors through an action of the quotient $\mathbb{T}^{F}\times^{\mathbb{Z}^{F}}\mathbb{T}^{F}$. This latter action extends to the action of $\mathbb{T}^{F}\times^{\mathbb{Z}^{F}}\mathbb{T}^{F}\subseteq(\mathbb{T}\times^{\mathbb{Z}}\mathbb{T})^{F}$ on $\Sigma^{i}$ (and $\Sigma^{i}_{w}$ for $w\in W_{e}$) given by the same formula. By the discussion in [DI24, §4.2] which applies in our more general setting, Theorem 1.1 follows from the next result.
Theorem 3.1.
Suppose that $q$ satisfies condition (2.1). Then there exists a Coxeter pair $(T,U)$ such that
$\dim_{\overline{}\mathbb{Q}_{\ell}}H_{\ast}(\Sigma_{w}^{0})=\begin{cases}H_{\ast}((\dot{w}\mathbb{T})^{c\sigma})&\text{if $w\in W_{e}^{c\sigma}$,}\\ 0&\text{otherwise.}\end{cases}$
as virtual $(\mathbb{T}\times^{\mathbb{Z}}\mathbb{T})^{F}$-modules.
As a first step towards the proof of Theorem 3.1 we observe that the whole discussion of [DI24, §4.3] applies mutatis mutandis in our setting. Thus it suffices to prove Theorem 3.1 in the case that $\Delta$ is connected. In particular, there exists some $c$ satisfying condition ($\ast$2.6), cf. Remark 2.4. Now Theorem 3.1 follows from Corollary 3.7 and Proposition 3.12 below.
3.2. An extension of action
Let $w\in W_{e}$. We set $K_{w,i}={}^{w^{-1}}U^{-}\cap F^{i}U^{-}$. Define
$\hat{\Sigma}_{w}^{i}=\{(\tilde{x},\tilde{x}^{\prime},y_{1},\tau,z,y_{2})\in F\mathbb{U}\times F^{i+1}\mathbb{U}\times\mathbb{U}\times{\dot{w}}\mathbb{T}\times\mathbb{K}_{w,i}^{1}\times F^{i}\mathbb{U};\tilde{x}F(\tau z)=y_{1}\tau zy_{2}\tilde{x}^{\prime}\}.$
We define an action of $\mathbb{T}^{F}\times\mathbb{T}^{F}$ on $\Sigma_{w}^{i}$ by
$(t,t^{\prime}):(\tilde{x},\tilde{x}^{\prime},y_{1},\tau,z,y_{2})(t\tilde{x}t^{-1},t^{\prime}\tilde{x}^{\prime}{t^{\prime}}^{-1},ty_{1}t^{-1},t\tau{t^{\prime}}^{-1},t^{\prime}z{t^{\prime}}^{-1},t^{\prime}y_{2}{t^{\prime}}^{-1}).$
Then there is an $\mathbb{T}^{F}\times\mathbb{T}^{F}$-equivariant affine space bundle
$\pi_{w}^{i}:\hat{\Sigma}_{w}^{i}\Sigma_{w}^{i},\quad(\tilde{x},\tilde{x}^{\prime},y_{1},\tau,z,y_{2})(\tilde{x}F(y_{1})^{-1},\tilde{x}^{\prime}F(y_{2}),y_{1}\tau zy_{2}).$
Let $\chi\in X_{*}(T)$ which centralizes $K_{w,i}$. Define
$H_{w,\chi}=\{(t,t^{\prime})\in\mathbb{T}_{h}\times\mathbb{T}_{h};w^{-1}t^{-1}F(t)w={t^{\prime}}^{-1}F(t^{\prime})\in\operatorname{Im}(\chi)\}.$
Then $H_{w,\chi}$ acts on $\hat{\Sigma}_{w}^{i}$ by
$(t,t^{\prime}):(\tilde{x},\tilde{x}^{\prime},y_{1},\tau,z,y_{2})({}^{F(t)}\tilde{x},{}^{F(t^{\prime})}\tilde{x}^{\prime},{}^{F(t)}y_{1},t\tau{t^{\prime}}^{-1},{}^{t^{\prime}}z,{}^{F(t^{\prime})}y_{2}).$
Lemma 3.2.
Let $w\in W_{e}\setminus W_{e}^{c\sigma}$ such that $\Sigma_{w}^{i}\neq\emptyset$. Then there exists a proper subset $K=\sigma_{I}(K)\subsetneq\Delta_{e}$ such that $w(c_{I}\sigma_{I})^{i}\sigma_{I}^{-i}\in w_{e}W_{K}$.
Proof.
Let $w_{i}=w(c_{I}\sigma_{I})^{i}\sigma_{I}^{-i}\in W_{e}$. By assumption we have
$c\sigma\mathbb{B}w(c\sigma)^{i}\mathbb{B}\mathbb{G}^{1}(c\sigma)^{-i-1}\cap\mathbb{B}w(c\sigma)^{i}\mathbb{B}\mathbb{G}^{1}c\sigma\mathbb{B}(c\sigma)^{-i-1}\neq\emptyset.$
As $c_{I}\sigma_{I}=c\sigma$, this implies that
$c_{I}\sigma_{I}\mathbb{B}_{1}w(c_{I}\sigma_{I})^{i}\mathbb{B}_{1}\cap\mathbb{B}_{1}w(c_{I}\sigma_{I})^{i}\mathbb{B}_{1}c_{I}\sigma_{I}\neq\emptyset,$
that is,
$c_{I}\mathbb{B}_{1}\sigma_{I}(w_{i})\mathbb{B}_{1}\cap\mathbb{B}_{1}w_{i}\mathbb{B}_{1}(\sigma_{I})^{i}(c_{I})\neq\emptyset.$
In particular there are $\sigma_{I}$-Coxeter elements $v^{\prime}\leq_{e}c_{I}$ and $v\leq_{e}(\sigma_{I})^{i}(c_{I})$ of some $\sigma_{I}$-stable subsets $K^{\prime}$ and $K$ of $\Delta_{e}$ respectively (one of which is a proper subset of $\Delta_{e}$ since $w\in W_{e}\setminus W_{e}^{c\sigma}$) such that $v^{\prime}\sigma_{I}(w_{i})=w_{i}v$ and
$\mathbb{B}_{1}w_{i}\mathbb{B}_{1}(\sigma_{I})^{i}(c_{I})\cap\mathbb{B}_{1}w_{i}v\mathbb{B}_{1}\neq\emptyset.$
Applying Lemma 2.7, there exist $x=\sigma(x)\in{}^{K^{\prime}}W_{e}{}^{K}$ such that $K^{\prime}={}^{x}K$ and $w_{i}\in xW_{K}$. Moreover, it follows from (a) that for any simple reflection $s\in\operatorname{supp}_{\Delta_{e}}(\sigma_{I}^{i}(c_{I}))\setminus K$ we have $xs\in W_{K^{\prime}}x$ or $xs\leq_{e}x$. The former is impossible since $s\notin W_{K}=xW_{K^{\prime}}x^{-1}$. So we have $xs\leq_{e}x$. Moreover, as $xsx^{-1}\notin W_{K^{\prime}}$ we have $w_{K^{\prime}}xs\leq_{e}w_{K^{\prime}}x=xw_{K}$, where $w_{K}$ and $w_{K^{\prime}}$ are the maximal elements of $W_{K}$ and $w_{K^{\prime}}$ respectively. As $xw_{K}$ is $\sigma_{I}$-stable, we have $xw_{K}s\leq_{e}xw_{K}$ for all $s\in\Delta_{e}$, that is, $xw_{K}=w_{e}$. Hence $w_{i}\in w_{e}W_{K}$. ∎
Let $N_{0}\in\mathbb{Z}_{\geqslant 0}$ be the order of $c\sigma\in W\rtimes\langle\sigma\rangle$. Define
$N_{F}^{F^{N_{0}}}:\mathbb{T}_{h}\mathbb{T}_{h},\quad ttF(t)\cdots F^{N_{0}-1}(t).$
Lemma 3.3.
Let $\chi\in X_{*}(T)$ and let $C$ be a $c\sigma$-orbit of $\Phi$. Assume $\chi$ is non-central on $C$ and $|\langle\chi,\beta\rangle|<q$ for $\beta\in C$. Then $\sum_{i=0}^{N_{0}-1}q^{i}\langle\gamma,(c\sigma)^{i}(\chi)\rangle\neq 0$ for $\gamma\in C$. In particular, the action of $\mathbb{G}_{m}$ on $\mathbb{U}_{\gamma}$ for $\gamma\in C$, via the morphism $N_{F}^{F^{N_{0}}}\circ\chi$, is nontrivial.
Proof.
By assumption, $|\langle\gamma,(c\sigma)^{i}(\chi)\rangle|=|\langle(c\sigma)^{-i}(\gamma),\chi\rangle|<q$ for $10\leqslant i\leqslant N_{0}-$ 1, and there exists $10\leqslant i_{0}\leqslant N_{0}-$ 1 such that $\langle(c\sigma)^{-i_{0}}(\gamma),\chi\rangle\neq 0$. Hence the statement follows. ∎
Let $\mathbb{G}_{m}\subseteq\mathcal{O}_{\breve{k}}^{\times}$ be the Teichmüller lift of the quotient map $\mathcal{O}_{\breve{k}}^{\times}\overline{}\mathbb{F}_{q}^{\times}$. Assume that $r\in\mathbb{Z}_{\geqslant 1}$.
Lemma 3.4.
Consider the homomorphism
$f_{w,\chi}:\mathbb{G}_{m}\mathbb{T}\times\mathbb{T},\quad x(N_{F}^{F^{N_{0}}}({}^{w}\chi(x)),N_{F}^{F^{N_{0}}}(\chi(x))).$
Then $\operatorname{Im}(f_{w,\chi})\subseteq H_{w,\chi}^{\circ}$.
Proof.
By definition. $F^{N_{0}}(\lambda(x))=\lambda(x^{q^{N_{0}}})$ for $x\in{\breve{k}}$. Hence
$N_{F}(\chi(x))^{-1}F(N_{F}(\chi(x)))=\chi(x)^{-1}F^{N_{0}}(\chi(x))=\chi(x^{-1}\sigma^{N_{0}}(x)).$
So the statement follows. ∎
3.3. Handling $\Sigma_{w}^{0}$ for $w\in W_{e}\setminus W_{e}^{c\sigma}$
Let $i\in\mathbb{Z}$. Following [DI24, §5] we define an isomorphism of varieties
$\alpha_{i}:\Sigma^{i}\Sigma^{i+1},\quad(x,x^{\prime},y)(x,F(x^{\prime}),yx^{\prime}).$
For $w,u\in W_{e}$ we define
$Y_{w,u}^{i}$ $=\Sigma_{w}^{i}\cap(\alpha_{i})^{-1}(\Sigma_{u}^{i+1});$
$Z_{w,u}^{i+1}$ $=\alpha_{i}(\Sigma_{w}^{i})\cap\Sigma_{u}^{i+1}=\alpha_{i}(Y_{w,u}^{i}).$
Let $\hat{Y}_{w,u}^{i}=(\pi_{w}^{i})^{-1}(Y_{w,u}^{i})$ and $\hat{Y}_{w,u}^{i}=(\pi_{u}^{i+1})^{-1}(Z_{w,u}^{i+1})$.
Lemma 3.5.
Let $w,u\in W_{e}$. Let $\chi,\mu\in X_{*}(T)$ which centralizes $\mathbb{K}_{w,i}$ and $\mathbb{K}_{u,i+1}$ respectively. Then $H_{w,\chi}$ and $H_{u,\mu}$ preserve $\hat{Y}_{w,u}^{i}$ and $\hat{Z}_{w,u}^{i+1}$ respectively.
Proof.
This is proved in [DI24, §5]. ∎
Proposition 3.6.
Suppose that condition (italic-$\ast$2.6) holds and that $q$ satisfies condition (2.1). Let $i\in\mathbb{Z}$. Then
$H_{*}(\hat{Y}_{w,u}^{i})=H_{*}(Y_{w,u}^{i})=H_{*}(Z_{w,u}^{i+1})=H_{*}(\hat{Z}_{w,u}^{i+1})=0$
if $w$ or $u$ belongs to $W_{e}\setminus W_{e}^{c\sigma}$.
Proof.
Without loss of generality we can assume that $w\in W_{e}\setminus W_{e}^{c\sigma}$ and $\hat{Y}_{w,u}^{i}\neq\emptyset$. In particular, $\Sigma_{w}^{i}\neq\emptyset$. By Lemma 3.2 and Corollary 2.6, there are subsets $K=\sigma_{I}(K)\subsetneq\Delta_{e}$ and $J=\sigma(J)\subsetneq\Delta$ such that
$w(c\sigma)^{i}\in w_{e}W_{K}(\sigma_{I})^{i}\subseteq w_{e}W_{J}\sigma^{i}=w_{0}W_{J}\sigma^{i}.$
Thus
$K_{w,i}\subseteq{}^{w^{-1}}(U^{-}\cap{}^{w(c\sigma)^{i}}U^{-})\subseteq{}^{w^{-1}w_{0}}M_{J},$
where $M_{J}$ is the Levi subgroup generated by $T$ and $U_{\gamma}$ for $\gamma\in\Phi_{J}$. Let $\mathcal{O}\in\Delta\setminus J$ be a $\sigma$-orbit. Then $W_{J}$ fixes $\omega_{\mathcal{O}}^{\vee}$, and $K_{w}\subseteq{}^{w^{-1}w_{0}}M_{J}$ is centralized by
$\chi:=w^{-1}w_{0}(\omega_{\mathcal{O}}^{\vee})=w^{-1}w(c\sigma)^{i}\sigma^{-i}(\omega_{\mathcal{O}}^{\vee})=(c\sigma)^{i}(\omega_{\mathcal{O}}^{\vee}).$
Moreover, $w(\chi)=w_{0}\sigma^{N}(\omega_{\mathcal{O}}^{\vee})=(c\sigma)^{N}(\omega_{\mathcal{O}}^{\vee})$.
Let $f_{w,\chi}:\mathbb{G}_{m}H_{w,\chi}$ be the as in Lemma 3.4. In view of Lemma 3.5, via $f_{w,\chi}$ the action of $H_{w,\chi}$ on $\hat{Y}_{w,u}^{i}$ induces an action of $\mathbb{G}_{m}$ on $\hat{Y}_{w,u}^{i}$, which commutes with action of $\mathbb{T}^{F}\times\mathbb{T}^{F}$. Hence
$H_{c}^{*}(Y_{w,u})=H_{c}^{*}(\hat{Y}_{w,u}^{i})=H_{c}^{*}((\hat{Y}_{w,u}^{i})^{\mathbb{G}_{m}}),$
it suffices to show $(\hat{Y}_{w,u}^{i})^{\mathbb{G}_{m}}=\emptyset$. To this end, we can assume that $\Delta=\cup_{i\in\mathbb{Z}}\sigma^{i}(H)$ for some/any connected component $H$ of $\Delta$. Then by Proposition 2.3, $\chi,w(\chi)\in\{(c\sigma)^{i}(\omega_{\mathcal{O}}^{\vee});i\in\mathbb{Z}\}$ are non-central on each $c\sigma$-orbit of $\Phi$. As $q>M$, it follows from Lemma 3.3 that
$(\hat{\Sigma}_{w,u}^{i})^{\mathbb{G}_{m}}\subseteq\{1\}\times\{1\}\times\{1\}\times\mathbb{T}\times\{1\}\times\{1\}.$
As $w\in W_{e}\setminus W_{e}^{c\sigma}$, we deduce that $(\hat{Y}_{w,u}^{i})^{\mathbb{G}_{m}}=\emptyset$ as desired. ∎
Corollary 3.7.
Let $i\in\mathbb{Z}$ and $w\in W_{e}$. If $w\in W_{e}\setminus W_{e}^{c\sigma}$ then $H_{*}(\Sigma_{w}^{i})=0$. Otherwise,
$H_{*}(\Sigma_{w}^{i})=\sum_{u\in W_{e}^{c\sigma}}H_{*}(Y_{w,u}^{i})=\sum_{u\in W_{e}^{c\sigma}}H_{*}(Z_{u,w}^{i})=\sum_{u\in W_{e}^{c\sigma}}H_{*}(Y_{u,w}^{i-1}).$
Proof.
Note that $\Sigma_{w}^{i}=\sqcup_{u\in W_{e}}Y_{w,u}^{i}=\sqcup_{u\in W_{e}}Z_{u,w}^{i}$ and $Z_{u,w}^{i}\cong Y_{u,w}^{i-1}$. Then the statement follows from Proposition 3.6. ∎
3.4. Handling $\Sigma_{w}^{0}$ for $w\in W_{e}^{c\sigma}$
Lemma 3.8.
Suppose that Condition (italic-$\ast$2.6) holds. Let $i\in\mathbb{Z}$ and $w,u\in W_{e}^{c\sigma}$ such that $Y_{w,u}^{i}\neq\emptyset$. Then $w=u$ if either $\sigma_{I}\neq 1$ or $\sigma_{I}=1$ and $wc_{I}^{i}\neq w_{e}$.
Proof.
By assumption, we have
$\mathbb{B}_{1}w(c_{I}\sigma_{I})^{i}\mathbb{B}_{1}c_{I}\sigma_{I}\mathbb{B}_{1}(c_{I}\sigma_{I})^{-i-1}\cap\mathbb{B}_{1}u(c_{I}\sigma_{I})^{i+1}\mathbb{B}_{1}(c_{I}\sigma_{I})^{-i-1}\neq\emptyset,$
that is, $\mathbb{B}_{1}w(c_{I}\sigma_{I})^{i}\mathbb{B}_{1}c_{I}\sigma_{I}\mathbb{B}_{1}\cap\mathbb{B}_{1}u(c_{I}\sigma_{I})^{i+1}\mathbb{B}_{1}\neq\emptyset$. Thus there exists $v\leq_{e}c_{I}$ such that $w(c_{I}\sigma_{I})^{i}v\sigma_{I}=u(c_{I}\sigma_{I})^{i+1}$. Note that $w,u\in W_{e}^{c\sigma}\subseteq\langle c_{I}\sigma_{I}\rangle$. We have
$v\sigma_{I}=(c_{I}\sigma_{I})^{-i}w^{-1}u(c_{I}\sigma_{I})^{i+1}=w^{-1}u(c_{I}\sigma_{I})\in\langle c_{I}\sigma_{I}\rangle.$
In particular, it follows from Lemma 2.5 that $\ell_{e}(v)$ is divided by $\ell_{e}(c_{I})$.
Assume that either $\sigma_{I}\neq 1$ or $\sigma_{I}=1$ and $wc_{I}^{i}\neq w_{e}$. If $v\neq 1$, then $\ell_{e}(v)=\ell_{e}(c_{I})$ since $1\neq v\leq c_{I}$. Hence we have $v=c_{I}$ and $w=u$ as desired. Suppose $v=1$. Then $c_{I}=u^{-1}w\in W_{e}^{c\sigma}$, which means that $\sigma_{I}(c_{I})=c_{I}$. Hence $\sigma_{I}=1$ by Theorem 2.1 (3). By assumption we have $\sigma_{I}=1$ and $w\sigma_{I}^{i}\neq w_{e}$. As $v=1$, we have $wc_{I}^{i}s<wc_{I}^{i}$ for all $s\in\operatorname{supp}_{\Delta_{e}}(c_{I})=\Delta_{e}$, that is, $wc_{I}^{i}=w_{e}$, a contradiction. ∎
Theorem 3.9 ([IN24], Theorem 3.1).
The map $(u_{1},u_{2})u_{1}^{-1}u_{2}F(u_{1})$ gives an isomorphism
$\phi:(F\mathbb{U}\cap\mathbb{U})\times(F\mathbb{U}\cap\mathbb{U}^{-})\cong F\mathbb{U}.$
In particular, $\phi$ restricts to an isomorphism
$(F\mathbb{U}^{1}\cap\mathbb{U})\times(F\mathbb{U}\cap\mathbb{U}^{-})\cong\mathbb{U}^{1}(F\mathbb{U}\cap\mathbb{U}^{-}).$
For $i\in\mathbb{Z}$ and $w\in W_{e}$ we define
${}^{\flat}\Sigma_{w}^{i}=\{(x,x^{\prime},y)\in(F\mathbb{U}\cap\mathbb{U}^{-})\times F^{i}(F\mathbb{U}\cap\mathbb{U}^{-})\times\mathbb{G};xF(y)=yx^{\prime}\}.$
Lemma 3.10.
The map $(x,x^{\prime},y)(x_{2},x_{2}^{\prime},x_{1}yF^{i}(x_{1}^{\prime})^{-1},x_{1},x_{1}^{\prime})$ gives an $\mathbb{T}^{F}\times\mathbb{T}^{F}$-equivariant isomorphism
$\Sigma_{w}^{i}\cong{}^{\flat}\Sigma_{w}^{i}\times(F\mathbb{U}\cap\mathbb{U})\times(F\mathbb{U}\cap\mathbb{U}),$
where $(x_{1},x_{2})=\phi^{-1}(x)$ and $(x_{1}^{\prime},x_{2}^{\prime})=\phi^{-1}(x^{\prime})$. In particular, $H_{c}^{*}(\Sigma_{w}^{i})\cong H_{c}^{*}({}^{\flat}\Sigma_{w}^{i})$.
Proof.
It follows by definition and Theorem 3.9. ∎
Lemma 3.11.
Suppose that $c$ satisfies condition (italic-$\ast$2.6). Let $w=(c\sigma)^{m}\in W_{e}^{c\sigma}$ for some $m\in\mathbb{Z}$. Then we have $H_{*}(\Sigma_{w}^{N-m})=H_{*}({\dot{w}}\mathbb{T}^{F})=H_{*}({}^{\flat}\Sigma_{w}^{2N-m})=H_{*}(\Sigma_{w}^{2N-m})$.
Proof.
The first statement is proved in [DI24]. We show the second one. Let $(x,x^{\prime},y)\in{}^{\flat}\Sigma_{w}^{-m}$. By definition,
$y\in\mathbb{G}^{1}\mathbb{B}{\dot{w}}(c\sigma)^{2N-m}\mathbb{B}(c\sigma)^{-m+2N}=\mathbb{U}\mathbb{T}\mathbb{U}^{-,1}{\dot{w}}.$
So we may write $y=y_{1}\tau y_{2}w$ uniquely with $y_{1}\in\mathbb{U}$, $\tau\in\mathbb{T}$ and $y_{2}\in\mathbb{U}^{-,1}$.
Then the equality $xF(y)=yx^{\prime}$ is equivalent to
$\tau^{-1}y_{1}^{-1}xF(y_{1})F(\tau)=y_{2}{\dot{w}}x^{\prime}{\dot{w}}^{-1}F(y_{2}^{-1})=y_{2}x^{\prime\prime}F(y_{2}^{-1}),$
where $x^{\prime\prime}={}^{\dot{w}}x^{\prime}\in F\mathbb{U}\cap\mathbb{U}^{-}$ since $w=(c\sigma)^{m}$.
By Theorem 3.9, the map $(g_{1},g_{2})g_{1}^{-1}g_{2}F(g_{1})$ gives isomorphisms
$\mathbb{U}\times(F\mathbb{U}\cap\mathbb{U}^{-})$ $\cong\mathbb{U}(F\mathbb{U}\cap\mathbb{U});$
$\mathbb{U}^{-,1}\times(F\mathbb{U}\cap\mathbb{U}^{-})$ $\cong(F\mathbb{U}\cap\mathbb{U}^{-})F\mathbb{U}^{-,1}.$
So we can make changes of variables $(x,x^{\prime\prime},y_{1},y_{2})(z_{1},z_{2},z_{3},z_{4})$, where
$(z_{1},z_{2},z_{3},z_{4})\in\mathbb{U}\times F\mathbb{U}\cap\mathbb{U}^{-}\times\mathbb{U}^{-,1}(F\mathbb{U}\cap\mathbb{U}^{-})\times F\mathbb{U}^{-,1}\cap\mathbb{U}$
such that $y_{1}^{-1}xF(y_{1})=z_{1}z_{2}$ and $y_{2}x^{\prime\prime}F(y_{2})^{-1}=z_{3}z_{4}$. Then we have
$\tau^{-1}z_{1}z_{2}F(\tau)={}^{\tau^{-1}}z_{1}L(\tau){}^{F(\tau)^{-1}}z_{2}=z_{3}z_{4},$
where $L(\tau)=\tau^{-1}F(\tau)$. As $z_{4}\in\mathbb{U}^{1}$ we can have
${}^{F(\tau)^{-1}}z_{2}z_{4}^{-1}=h_{+}h_{0}h_{-}\in\mathbb{U}\mathbb{T}\mathbb{U}^{-},$
where $h_{+}\in\mathbb{U}^{1}$, $h_{0}\in\mathbb{T}^{1}$ and $h_{-}\in(F\mathbb{U}\cap\mathbb{U}^{-})\mathbb{U}^{-,1}=F(\mathbb{U}\mathbb{U}^{-,1})\cap\mathbb{U}^{-}$. Hence
${}^{\tau^{-1}}z_{1}{}^{L(\tau)}h_{+}L(\tau)h_{0}h_{-}=z_{3}.$
It follows that $z_{1}={}^{F(\tau)}h_{+}$, $L(\tau)=h_{0}^{-1}$ and $z_{3}=h_{-}$. Therefor,
$\Sigma_{w}^{i}=\{(\tau,z_{2},z_{4})\in\mathbb{T}\times(F\mathbb{U}\cap\mathbb{U}^{-})\times(F\mathbb{U}^{-,1}\cap\mathbb{U});L(\tau)={\rm pr}_{0}({}^{F(\tau)^{-1}}z_{2}z_{4})\},$
where ${\rm pr}_{0}:\mathbb{U}^{1}\mathbb{T}\mathbb{U}^{-}\mathbb{T}$ is the natural projection.
Note that $(t,t^{\prime})\in\mathbb{T}^{F}\times\mathbb{T}^{F}$ acts on ${}^{\flat}\Sigma_{w}^{i}$ by $(\tau,z_{2},z_{4})(t\tau w(t^{\prime})^{-1},{}^{t}z_{2},{}^{w(t^{\prime})}z_{4})$. Now we define and action of $s\in\mathbb{T}$ on ${}^{\flat}\Sigma_{w}^{i}$ by $(\tau,z_{2},z_{4})(\tau,{}^{s}z_{2},{}^{s}z_{4})$. Then the actions of $\mathbb{T}$ and $\mathbb{T}^{F}\times\mathbb{T}^{F}$ commutes with each other. Thus, by Lemma 3.10 we have
$H_{*}(\Sigma_{w}^{2N-m})=H_{*}({}^{\flat}\Sigma_{w}^{2N-m})\cong H_{*}((\Sigma_{w}^{i})^{\mathbb{T}})=H_{*}({\dot{w}}\mathbb{T}^{F})$
as desired. ∎
Proposition 3.12.
Suppose that Condition (italic-$\ast$2.6) holds and that $\Delta$ is connected. Then $H_{c}^{*}(\Sigma_{w}^{0})=H_{c}^{*}({\dot{w}}\mathbb{T}^{F})$ for $w\in W_{e}^{c\sigma}$.
Proof.
Let $w\in W_{e}^{c\sigma}$. As $\Delta$ is connected, we may write $w=(c\sigma)^{m}$ for some $m\in\mathbb{Z}$. By Corollary 3.7 we have
$H_{*}(\Sigma_{w}^{i})=\sum_{u\in W_{e}^{c\sigma}}H_{*}(Y_{w,u}^{i}),\quad H_{*}(\Sigma_{w}^{i+1})=\sum_{u\in W_{e}^{c\sigma}}H_{*}(Y_{u,w}^{i}).$
First we assume $\sigma_{I}\neq 1$. By Lemma 3.8 for any $w^{\prime},u^{\prime}\in W_{e}^{c\sigma}$ we have $Y_{w^{\prime},u^{\prime}}^{i}\neq\emptyset$ if and only if $w^{\prime}=u^{\prime}$. It follows by (a) that
$H_{*}(\Sigma_{w}^{i})=H_{*}(Y_{w,w}^{i})=H_{*}(\Sigma_{w}^{i+1}).$
By Lemma 3.11 we have $H_{*}(\Sigma_{w}^{0})=H_{*}(\Sigma_{w}^{N-m})=H_{*}(w\mathbb{T}^{F})$ as desired.
Now we assume $\sigma_{I}=1$. Let notation be as in Lemma 2.5. We can assume that $w=c_{I}^{m}$ with $10\leqslant m\leqslant 2N-$ 1. If $0\leqslant m\leqslant N$, it follows from (a), Lemma 3.8 and Lemma 3.11 that
$H_{*}(\Sigma_{w}^{0})=H_{*}(\Sigma_{w}^{1})=\cdots=H_{*}(\Sigma_{w}^{N-m})=H_{*}({\dot{w}}\mathbb{T}^{F}).$
If $N+1\leqslant m\leqslant 2N-1$, similarly we have
$H_{*}(\Sigma_{w}^{0})=H_{*}(\Sigma_{w}^{1})=\cdots=H_{*}(\Sigma_{w}^{2N-m})=H_{*}({\dot{w}}\mathbb{T}^{F}).$
So the statement follows. ∎
4. Proof of Theorem 2.1
In this section, we fill in the proof for Theorem 2.1. First we show that it suffices to consider one particular Coxeter element.
Lemma 4.1.
Let $\alpha\in\{\alpha_{1},\sigma^{-1}(\alpha_{r})\}$ such that $c^{\prime}=s_{\alpha}c\sigma(s_{\alpha})$. Suppose Theorem 2.1 holds for $(\mu,c)$. Then it also holds for $(s_{\alpha}(\mu),c^{\prime})$.
Proof.
Let $\mu,e,I$ be as in Theorem 2.1. Let $e^{\prime}=e_{s_{\alpha}(\mu),c^{\prime}}=s_{\alpha}(e)$ and $\Phi_{e^{\prime}}=s_{\alpha}(\Phi_{e})$. Assume that $I=(i_{1}<\cdots<i_{m})$. Without loss of generality we can assume $\alpha=\sigma^{-1}(\alpha_{r})$ and $c^{\prime}=s_{\alpha_{1}^{\prime}}s_{\alpha_{2}^{\prime}}\cdots s_{\alpha_{r}^{\prime}}$ with $\alpha_{1}^{\prime}=\alpha_{r}$ and $\alpha_{i}^{\prime}=\alpha_{i-1}$ for $2\leqslant i\leqslant r$.
First we assume that $r\in I$. Then $r=i_{m}$ and $\sigma_{I,c}(\alpha)<0$, which means that $\alpha\notin\Delta_{e}=\sigma_{I,c}(\Delta_{e})$. Thus $\Phi_{e^{\prime}}^{+}=s_{\alpha}(\Phi_{e}^{+})$ since $\alpha\in\Delta$ is a simple root. In particular, $\Delta_{e^{\prime}}=s_{\alpha}(\Delta_{e})$. We take
$I^{\prime}=(1<i_{1}+1<i_{2}+1<\cdots<i_{m-1}+1).$
Then $\sigma_{I^{\prime},c^{\prime}}=s_{\alpha}\sigma_{I,c}s_{\alpha}$, $c^{\prime}_{I^{\prime}}=s_{\alpha}c_{I}s_{\alpha}$ and the statement follows.
Now we assume that $r\notin I$. Then $\sigma_{I,c}(\alpha)\in\Delta_{I,c}\subseteq\Delta_{e}=\sigma_{I,c}(\Delta_{e})$. Thus $\alpha\in\Delta_{e}$ and $\Delta_{e^{\prime}}=\Delta_{e}$. We take
$I^{\prime}=(i_{1}+1<i_{2}+1<\cdots<i_{m}+1).$
Then $\sigma_{I^{\prime},c^{\prime}}=\sigma_{I,c}$, $c_{I^{\prime}}^{\prime}=s_{\alpha}c_{I}\sigma_{I,c}(s_{\alpha})$ and the statement also follows. ∎
To finish the proof, we will take a particular $\sigma$-Coxeter element $c$ such that and verify the statement directly. Moreover, we can assume $\Delta$ is connected.
Let $P$ be the coweight be the coweight lattice of $\Phi$. If $\mu=0\in P/(1-c\sigma)P$, then $\Delta_{e}=\Delta$ and the statement is trivial. So we may assume that $X_{*}(T)/(1-c)X_{*}(T)\neq\{0\}$, which excludes the types ${}^{2}A_{n-1}$ ($n$ even), ${}^{3}D_{4}$, $E_{8}$, ${}^{2}E_{6}$, $F_{4}$, $G_{2}$. Then we will take a case-by-case analysis for the remaining types.
We adopt the labeling of Dynkin diagrams by positive integers as in [Hum72]. For $i\in\mathbb{Z}_{\geqslant 1}$. let $s_{i}$ and $\omega_{i}^{\vee}$ denote the corresponding simple reflection and fundamental coweight, respectively.
Case (1): $\Delta$ is of type $A_{n-1}$. Take $c=s_{1}s_{2}\cdots s_{n-1}$. Then we have $P/(1-c\sigma)P=\{0,\omega_{1}^{\vee},\omega_{2}^{\vee},\dots\omega_{n-1}^{\vee}\}$. Assume $\mu=\omega_{k}^{\vee}$ with $k\in\mathbb{Z}$. Let $m={\rm{gcd}}(k,n)\in\mathbb{Z}_{\geqslant 1}$. Then we take $I$ to be the complement of the sequence $I^{c}=(n/m,2n/m,\cdots,(m-1)n/m)$.
Case (2): $\Delta$ is of type ${}^{2}A_{n-1}$ with $n\geqslant 4$ even. Take $c=s_{1}s_{2}\cdots s_{n/2}$. Then $P/(1-c\sigma)P=\{0,\omega_{1}^{\vee}\}$. Assume $\mu=\omega_{1}^{\vee}$. Then we take $I=(n/2)$.
Case (3): $\Delta$ is of type $B_{n}$ with $n\geqslant 2$. Take $c=s_{1}s_{2}\cdots s_{n}$. Then $P/(1-c\sigma)P=\{0,\omega_{1}^{\vee}\}$. Assume $\mu=\omega_{1}^{\vee}$. Then we take $I=(n)$.
Case (4): $\Delta$ is of type $C_{n}$ with $n\geqslant 3$. Take $c=s_{1}s_{2}\cdots s_{n}$. Then $P/(1-c\sigma)P=\{0,\omega_{n}^{\vee}\}$. Assume $\mu=\omega_{n}^{\vee}$. Then we take $I=(1,3,\dots,n-\frac{(-1)^{n}+1}{2})$.
Case (5): $\Delta$ is of type $D_{n}$ with $n\geqslant 4$. Take $c=s_{1}s_{2}\cdots s_{n}$. Then $P/(1-c\sigma)P=\{0,\omega_{1}^{\vee},\omega_{n-1}^{\vee},\omega_{n}^{\vee}\}$. If $\mu=\omega_{1}^{\vee}$, take $I=(n-1,n)$. It remains to handle the case $\mu=\omega_{n-1}^{\vee}$ by symmetry. If $n$ is even, take $I=(1,3,\dots,n-3,4)$ if $4\mid n$ and $I=(1,3,\dots,n-3,n-1)$ if $4\nmid n$. If $n$ is odd, take $I=(1,3,\dots,n-3,n-4,\dots,n-2,n)$.
Case (6): $\Delta$ is of type $E_{6}$. Take $c=s_{1}s_{3}s_{4}s_{2}s_{5}s_{6}$. Then $P/(1-c\sigma)P=\{0,\omega_{1}^{\vee},\omega_{6}^{\vee}\}$. By symmetry we can assume $\mu=\omega_{1}^{\vee}$. Then take $I=(1,3,5,6)$.
Case (7): $\Delta$ is of type $E_{6}$. Take $c=s_{7}s_{6}s_{5}s_{4}s_{2}s_{3}s_{1}$. Then $P/(1-c\sigma)P=\{0,\omega_{7}^{\vee}\}$. If $\mu=\omega_{7}^{\vee}$, take $I=(7,5,2)$.